If a really large quantity of aqueous sodium chloride solution is reacted with 1.00 Kg of lead (II) Nitrate, how many grams of sodium chloride will be used in the reaction?
1
Expert's answer
2018-02-22T06:36:52-0500
Solution
Write the reaction equation:
Pb(NO3)2 + 2NaCl  PbCl2 + 2NaNO3
So 1 mole of Pb(NO3)2 reacts with 2 moles of NaCl.
Find the amount of Pb(NO3)2 (331 g/mole):
ν =
1000
331
= 3.02 (mole)
3.02 moles of Pb(NO3)2 react with 6.04 moles of NaCl.
Find the mass of NaCl:
m = 6.04 × 58.5 = 353.3 (g)
Answer
353.3 g of sodium chloride will be used in the reaction.